Compute Lead Time Williams Optical Inc. is considering a new lean product cell. The present manufacturing...

  1. Compute Lead Time

    Williams Optical Inc. is considering a new lean product cell. The present manufacturing approach produces a product in four separate steps. The production batch sizes are 45 units. The process time for each step is as follows:

    Process Step 1 5 minutes
    Process Step 2 8 minutes
    Process Step 3 4 minutes
    Process Step 4 3 minutes

    The time required to move each batch between steps is 5 minutes. In addition, the time to move raw materials to Process Step 1 is also 5 minutes, and the time to move completed units from Process Step 4 to finished goods inventory is 5 minutes.

    The new lean layout will allow the company to reduce the batch sizes from 45 units to 3 units. The time required to move each batch between steps and the inventory locations will be reduced to 2 minutes. The processing time in each step will stay the same.

    Determine the value-added, non-value-added, and total lead times, and the value-added ratio under the present and proposed production approaches. If required, round percentages to one decimal place.

    Present Approach Proposed Approach
    Value-added time min min
    Non-value-added time min min
    Total lead time min min
    Value-added ratio (as a percent) % %
Answer #1

Present Approach

Proposed Approach

Value-added time

20

min

20

min

Non-value-added time

905

min

50

min

Total lead time

925

min

70

min

Value-added ratio (as a percent)

2.2

%

28.6

%

Present Approach :-

Value added time = Total time of all steps

= 5 + 8 + 4 + 3 = 20 min

Non-value-added time = With in batch wait time + Move Time

With in batch wait time = Value added time * Remaining no of units after first unit start

                 = 20 min * 44 = 880 min

Move Time = 5 * 5 min = 25 min

Non-value-added time = 880 + 25 = 905 min

Value-added ratio (as a percent) = Value added time/Total lead time * 100

            = 20/925 * 100 = 2.2 %

Proposed Approach :-

Value added time = Total time of all steps

= 5 + 8 + 4 + 3 = 20 min

Non-value-added time = With in batch wait time + Move Time

With in batch wait time = Value added time * Remaining no of units after first unit start

                 = 20 min * 2 = 40 min

Move Time = 5 * 2 min = 10 min

Non-value-added time = 40 + 10 = 50 min

Value-added ratio (as a percent) = Value added time/Total lead time * 100

            = 20/70 * 100 = 28.6 %